OpenStudy (anonymous):

Solve the inequality: 7x-12<=to 9(2x-3)

OpenStudy (anonymous):

\[7x-12\leq 9(2x-3)\] \[7x-12\leq 18x-27\] \[-12\leq 11x-27\] \[15\leq 11x\] \[\frac{15}{11}\leq x\]